# Conformal Prediction: A Data Perspective

### v2

@@ Abstract.
Conformal prediction (CP), a distribution-free uncertainty quantification (UQ) framework, reliably provides valid predictive inference for black-box models. CP constructs prediction sets that contain the true output with a specified probability. However, modern data science’s diverse modalities, along with increasing data and model complexity, challenge traditional CP methods. These developments have spurred novel approaches to address evolving scenarios. This survey reviews the foundational concepts of CP and recent advancements from a data-centric perspective, including applications to structured, unstructured, and dynamic data. We also discuss the challenges and opportunities CP faces in large-scale data and models.
@@ 1. Introduction
The recent rapid development of well-designed and powerful machine learning (ML) models has significantly transformed our lives. However, the success of these models is often evaluated based on the accuracy of their predictions, which, while important, is not sufficient in many real-world scenarios. In high-stakes applications, it is equally critical to assess the uncertainty of model outputs. Uncertainty quantification (UQ) has long been a central problem in fields like statistics and ML. Several well-established methods, such as Bayesian inference and resampling techniques, have been widely adopted to address UQ. However, Bayesian posterior intervals are only valid if the parametric assumptions of the model are correctly specified, which may not always be the case in practical applications. Similarly, when constructing prediction sets for unobserved outputs, resampling approaches based on the residuals of observed data can result in overly narrow intervals due to overfitting, compromising the validity of the prediction set. This raises a fundamental question:
@@
CP (Vovk et al., 2005) is model-agnostic and can be applied to any ML model, generating prediction sets/intervals based on non-conformity scores derived from the model’s outputs. CP offers several desirable features: (1) Rigor/Finite-sample validity: CP generates calibrated prediction sets/intervals that ensure high-probability coverage with theoretical guarantees, without simplifying the models’ complexity or relying on strong (parametric) assumptions about the data generating process. More concretely, the validity of CP only requires the exchangeability assumption of data. (2) Lightweight: As a wrapper method, split CP, a commonly adopted variant, does not require retraining the underlying model, making it computationally efficient and easy to apply to existing models. (3) Versatility: CP is model-agnostic, meaning it can be used with any ML model, regardless of its architecture or complexity. (4) Informativeness: When appropriately designed, CP’s prediction set can be sufficiently small to provide meaningful information about the original outputs.
@@ 1.1. CP from a Data Perspective
The increasing prevalence of big data, particularly with richer and more complex modalities, presents unique challenges for CP beyond traditional statistical settings. For instance: (1) Generative models in natural language and image processing: In these contexts, the target output can be ultra high-dimensional, such as a paragraph or an image, unlike the univariate real-valued responses in classical regression or classification tasks. For complex tasks like open-domain question answering, the output space often has an infinite dimensionality (i.e., all possible text sequences). Adapting CP to handle these multi-dimensional and even infinite-dimensional outputs is essential to improve its applicability and predictive accuracy. (2) Time series and streaming data: These types of data are becoming increasingly important in fields such as meteorology, geosciences, urban planning, and healthcare. However, time-dependent data poses significant challenges for CP, particularly because the assumption of data exchangeability is often violated in the presence of distribution shifts or temporal dependencies. (3) Structured data (e.g., groups, trees, graphs): In machine learning applications involving structured data, canonical CP methods may fail to capture the intrinsic dependencies when the feature structure is ignored. Traditional inferential methods often rely on distributional assumptions specific to structured data. Thus, developing model-agnostic UQ tools that leverage implicit (or weighted) exchangeability within the data structure is crucial.
These challenges motivate this survey on CP from a data perspective. In this survey, we introduce CP methods tailored to different data categories, as outlined in Figure 1. We categorize data into two main types based on the collection process: static data and dynamic data. Static data is further divided into structured data—including matrix/tensor data, hierarchical data, tree-structured data, and graph data—and unstructured data, such as image, text, and multi-modal data. Dynamic data is split into streaming data and spatio-temporal data, with spatio-temporal data further subdivided into one-dimensional and multi-dimensional categories. This hierarchical organization provides a clear and systematic framework for understanding various data types, allowing for the appropriate application of CP techniques.

### v3

@@ Abstract.
Conformal prediction (CP), a distribution-free uncertainty quantification (UQ) framework, reliably provides valid predictive inference for black-box models. CP constructs prediction sets or intervals that contain the true output with a specified probability. However, modern data science’s diverse modalities, along with increasing data and model complexity, challenge traditional CP methods. These developments have spurred novel approaches to address evolving scenarios. This survey reviews the foundational concepts of CP and recent advancements from a data-centric perspective, including applications to structured, unstructured, and dynamic data. We also discuss the challenges and opportunities CP faces in large-scale data and models.
@@ 1. Introduction
The recent rapid development of well-designed and powerful machine learning (ML) models has significantly transformed our lives. However, the success of these models is often evaluated based on the accuracy of their predictions, which, while important, is not sufficient in many real-world scenarios. In high-stakes applications, it is equally critical to assess the uncertainty of model outputs. Uncertainty quantification (UQ) has long been a central problem in fields like statistics and ML. Several well-established methods, such as Bayesian inference and resampling techniques, have been widely adopted to address UQ. However, Bayesian posterior intervals are only valid if the distributional assumptions of the model are correctly specified, which may not always be the case in practical applications. Similarly, when constructing prediction sets for unobserved outputs, resampling approaches based on the residuals of observed data can result in overly narrow intervals since the estimate of variance is too small (Bates et al., 2024), compromising the validity of the prediction set. This raises a fundamental question:
@@
CP (Vovk et al., 2005) is model-agnostic and can be applied to any ML model, generating prediction sets/intervals based on non-conformity scores derived from the model’s outputs. CP offers several desirable features: (1) Finite-sample validity: CP generates calibrated prediction sets/intervals that ensure high-probability coverage with finite-sample theoretical guarantees, without simplifying the models’ complexity or relying on strong assumptions about the data generating process. More concretely, the validity of CP only requires the exchangeability assumption of data. (2) Lightweight: As a wrapper method, CP does not require retraining the underlying model, making it computationally efficient and easy to apply to existing models. (3) Versatility: CP is model-agnostic, meaning it can be used with any model, regardless of its architecture or complexity. (4) Informativeness: When appropriately designed, CP’s prediction set or interval can be sufficiently small to provide meaningful information about the original outputs.
@@ 1.1. CP from a Data Perspective
The increasing prevalence of big data, particularly with richer and more complex modalities, presents unique challenges for CP beyond traditional statistical settings. For instance: (1) Generative models in natural language and image processing: In these contexts, the target output can be ultra-high-dimensional or even infinite. For example, the output of a language model can be a permutation of any tokens in the pre-training corpus with arbitrary length , unlike the univariate real-valued responses in classical regression or classification tasks. For complex tasks like open-domain question answering, the output space is often unbounded (Su et al., 2024; Fisch et al., 2021b) (i.e., all possible text sequences). Adapting CP to handle these multi-dimensional and even infinite-dimensional outputs is essential to improve its applicability and predictive accuracy. (2) Time series and streaming data: These types of data are becoming increasingly important in fields such as economics, meteorology, geosciences, urban planning, and healthcare. However, time-dependent data poses significant challenges for CP, particularly because the assumption of data exchangeability is often violated in the presence of distribution shifts or temporal dependencies. (3) Structured data (e.g., groups, trees, graphs): Structured data is typically quantitative and organized in predefined formats like rows and columns in relational databases while unstructured data (like text) is often qualitative and lacking a fixed schema (Zhang et al., 2020; Tang et al., 2014). In ML applications involving structured data, canonical CP methods may fail to capture the intrinsic dependencies when the feature structure is ignored. Traditional inferential methods often rely on distributional assumptions specific to structured data. Thus, developing model-agnostic UQ tools that leverage implicit (or weighted) exchangeability within the data structure is crucial.
These challenges motivate this survey on CP from a data perspective, which introduces CP methods tailored to different data categories, as outlined in Figure 1. We categorize data into two main types based on the collection process: static data and dynamic data. Static data is further divided into structured data—including matrix/tensor data, hierarchical data, tree-structured data, and graph data—and unstructured data including image, text, and multi-modal data. Dynamic data is focused on spatio-temporal data, which is further subdivided into streaming, one-dimensional, and multi-dimensional categories. Note that dynamic data can be also structured, e.g., time-series data. This hierarchical organization provides a clear and systematic framework for understanding various data types, allowing for the appropriate application of CP techniques.